# Unveiling AI’s Blind Spots: An Oracle for In-Domain, Out-of-Domain, and Adversarial Errors

### v1

@@ Abstract
AI models make mistakes when recognizing images—whether in-domain, out-of-domain, or adversarial. Predicting these errors is critical for improving system reliability, reducing costly mistakes, and enabling proactive corrections in real-world applications such as healthcare, finance, and autonomous systems. However, understanding what mistakes AI models make, why they occur, and how to predict them remains an open challenge. Here, we conduct comprehensive empirical evaluations using a "mentor" model —a deep neural network designed to predict another model’s errors. Our findings show that the mentor model excels at learning from a mentee’s mistakes on adversarial images with small perturbations and generalizes effectively to predict in-domain and out-of-domain errors of the mentee. Additionally, transformer-based mentor models excel at predicting errors across various mentee architectures. Subsequently, we draw insights from these observations and develop an "oracle" mentor model, dubbed SuperMentor, that achieves 78% accuracy in predicting errors across different error types. Our error prediction framework paves the way for future research on anticipating and correcting AI model behaviours, ultimately increasing trust in AI systems. All code, models, and data will be made publicly available.
@@ 1 Introduction
Accurately predicting these errors is critical to enhancing the overall robustness and reliability of AI systems, especially in high-stakes real-world applications such as healthcare (Habehh & Gohel, 2021), finance (Mashrur et al., 2020), and autonomous driving (Huang et al., 2022). Proactively identifying potential errors enables more efficient corrections, reducing costly mistakes and safeguarding against catastrophic failures. By predicting when models are likely to err, we can implement strategies that either mitigate or entirely avoid the risks associated with those errors, ultimately leading to more trustworthy AI deployments.
@@
Subsequent research in out-of-domain detection has demonstrated that a model’s accuracy is often correlated with how far the data deviates from in-domain samples (Hendrycks & Dietterich, 2019; Shankar et al., 2021; Li et al., 2017). These methods typically rely on predefined metrics, such as model parameter distances (Yu et al., 2022), model disagreements (Jiang et al., 2021; Madani et al., 2004) and confidence scores (Guillory et al., 2021), which limits their ability to generalize predictions across various data types, including errors arising from in-domain data or adversarial attacks (Szegedy, 2013). Another line of research improves the robustness of the AI models with adversarial training approaches(Ilyas et al., 2019; Gowal et al., 2020; Balunović & Vechev, 2020); however, these approaches primarily focus on improving the model’s overall performance rather than predicting when errors may occur in the models.
@@
Different from all these previous works, we delve into the underlying principles of errors generated by AI models in the task of image classification with another AI model. Specifically, we designate the AI model that predicts errors as the mentor and the AI model being evaluated for performance as the mentee. The mentor strives to predict whether the mentee makes a mistake for any given test data. See Fig. 1 for the detailed illustration of the problem setup. Training the mentor on the error patterns made by the mentee can potentially reveal the strengths and weaknesses of the mentee’s learned representations across various visual contexts.
Specifically, we examine the effects of three distinct error types AI models often make: In-Domain (ID) Errors, Out-of-Domain (OOD) Errors, and Adversarial Attack (AA) Errors on three increasingly complex image datasets CIFAR-10 (Krizhevsky et al., 2009), CIFAR-100 (Krizhevsky et al., 2009) and ImageNet-1K (Deng et al., 2009). We identify which of these error types has the most significant impact on the mentor’s error prediction performances, and explore the reasons behind its prominence. Additionally, we assess how different mentor architectures influence error prediction accuracy and evaluate the mentor’s generalization performance across various mentee architectures. Finally, we develop a SuperMentor model that successfully predicts errors of the mentee with 78% accuracy across diverse error types. Our main contributions are highlighted below:
@@
1. We conduct an in-depth analysis of how training mentors on each of three distinct error types specified by the mentees—In-Domain (ID) Errors, Out-of-Domain (OOD) Errors, and Adversarial Attack (AA) Errors—affect the performance of error predictions over three increasingly complex image datasets. Our results reveal that training mentors with adversarial attack errors from the mentee has the most significant impact on improving the mentor’s error prediction accuracy.
@@
2. We explore how various mentor model architectures affect error prediction performance. Our experiments demonstrate that transformer-based mentor models outperform other architectures in accurately predicting errors.
@@
3. We investigate how varying levels of distortion in OOD and adversarial images affect the accuracy of error predictions. The findings indicate that training mentors with images with small perturbations can improve error prediction accuracy. In addition, we show that a mentor trained to learn error patterns from one mentee can successfully generalize its error predictions to another mentee.
@@ 2 Related Work
With the growing deployment of AI models across diverse fields, ensuring their reliability and understanding their limitations has become increasingly crucial. This has led to numerous research in safe AI such as uncertainty estimation (Nado et al., 2021; Lakshminarayanan et al., 2017), anomaly detection (Bogdoll et al., 2022), outlier detection (Boukerche et al., 2020), and out-of-domain detection (Yang et al., 2024). Unlike these areas, which mainly aim to predict whether the input data falls outside the training domain, our focus is on monitoring and predicting errors in AI models by determining whether the model’s output is correct, irrespective of whether the data comes from the training domain.
Moreover, to detect whether the input data is out of scope, the prior approaches mainly rely on softmax outputs (Granese et al., 2021; Hendrycks & Gimpel, 2016; Dang et al., 2024) or activations from network layers (Wang et al., 2020; Cheng et al., 2019; Ferreira et al., 2023), in applications such as object detection (Kang et al., 2018) and trajectory prediction (Shao et al., 2023; 2024). However, these methods often depend on manually defined metrics to estimate the probability of a mentee making a mistake. In contrast, our approach leverages another AI model to automatically learn and approximate the mentee’s decision boundaries, predicting its errors in an end-to-end trainable manner.
@@
Recent studies (Hendrycks & Dietterich, 2019; Shankar et al., 2021; Li et al., 2017) have shown that a model’s accuracy on a given dataset is often correlated with how far the data deviates from in-domain samples. However, these studies typically rely on pre-defined metrics, such as model parameter distances (Yu et al., 2022), model disagreements (Jiang et al., 2021; Madani et al., 2004), confidence scores (Guillory et al., 2021), domain-invariant representations (Chuang et al., 2020), and domain augmentation (Deng et al., 2021a), limiting their ability to generalize error prediction for in-domain data. In contrast, our mentor model is capable of predicting both OOD and in-domain errors for a mentee. Additionally, our mentor is an AI model trained end-to-end, eliminating the need for manually defined criteria.
@@ 3 Experimental setups
We denote the mentor and mentee networks as and respectively. We also define as the domain-specific set containing all the test images for a mentee, and as their ground-truth object class labels. Therefore, a mentor is expected to make perfect predictions about the correctness of the mentee’s responses (1 for "correct" and 0 for "wrong") given any image from :
@@ 3.1 Mentors
The first stream generates logits by predicting the probability distribution of a mentee over all the object classes when the mentee classifies the given image. The mentee network is kept fixed while training the mentor. Let us define the mentee’s output logit as . We introduce the distillation loss proposed by (Hinton, 2015): to align with . We set the temperature hyper-parameter in as 1.0, which controls the smoothness of the soft probability distribution. Higher temperatures make the distribution softer and more uniform across classes.
In the second stream, the mentor is prompted to predict whether the mentee will make a mistake on the given image or not. We denote the predicted binary label as , where 1 indicates that the mentee does not make a mistake and vice versa for 0. This prediction is supervised by where is the logistic regression loss and is the ground truth correctness label of a mentee. The overall loss is .
@@ 3.2 Mentees and their datasets
We employ two architectures as the mentees’ backbones: ResNet50 (He et al., 2016), which is a 2D-Convolutional neural network (2D-CNN), and ViT (Dosovitskiy, 2020), which is a transformer architecture based on self-attention mechanisms.
To train and test our mentees, we include three prominent image datasets of varying sizes and follow their standard data splits: CIFAR-10 (C10, (Krizhevsky et al., 2009)) with 10 object classes, CIFAR-100 with 100 object classes (C100, (Krizhevsky et al., 2009)) and ImageNet-1K with 1000 object classes (IN, (Deng et al., 2009)). Their multi-class recognition accuracy on the standard test sets of C10, C100 and IN datasets are 96.98%, 84.54%, 76.13% for the ResNet50 mentee and 97.45%, 86.51%, 81.07% for the ViT mentee respectively. The parameters of the mentees are frozen throughout all the experiments conducted on mentors.
@@ 3.3 Datasets for training and testing mentors
The mentor’s objective is to predict whether the mentee will misclassify a given image, regardless of its source. The mentor is trained on correctly and wrongly classified images by a mentee. Next, we introduce how these images are curated and collected.
A mentee may encounter various types of errors when dealing with real-world data. To explore which error types most effectively reveal the mentee’s learning patterns, we categorize errors into three types: (1) errors from in-domain test images, (2) errors from out-of-domain images, and (3) errors from adversarial images generated using adversarial attack methods. Next, we introduce these three error types in detail.
In-Domain (ID) Errors. occur on data that come from the same domain as the mentee’s training dataset. Specifically, errors on images from the standard validation set of ImageNet-1K or the test sets of CIFAR-10 and CIFAR-100 are considered ID errors. Along with the correctly classified images from these standard test sets, we create three datasets for a mentor: IN-ID, C10-ID, and C100-ID, following the naming convention of [Dataset]-[Error Type].
Out-of-domain (OOD) Errors. refer to errors that arise when the mentee encounters data outside the training domain. To obtain OOD samples of a dataset, we adopt four types of image corruptions from (Hendrycks & Dietterich, 2019): speckle noise (SpN) (noise category), Gaussian blur (GaB) (blur category), spatter (Spat) (weather category), and saturate (Sat) (digital category). The noise levels can vary and we select level 1 for image corruptions as specified in (Hendrycks & Dietterich, 2019) by default. As noise levels increase, the distortions on OOD images become more pronounced, leading to more mistakes of a mentee.
@@
Note that adversarial attacks are not always successful, and mentees can still correctly classify some adversarial images. We collect both the correctly and incorrectly classified adversarial images by a mentee based on C10 images, curating four datasets for the mentor: C10-AA-PGD, C10-AA-CW, C10-AA-Jitter, C10-AA-PIFGSM. Without the loss of generality, we can also curate four datasets each for a mentor based on C100 and IN images of a mentee.
Training and Test Splits.
@@
For any given dataset of a mentor, let and represent the sets of correctly and incorrectly classified images by a mentee. The sizes of and can vary significantly, depending on the mentee’s classification performance. A mentee with high recognition accuracy will have more correct classifications (big ) and fewer incorrect ones (small ). To create a balanced test set for a mentor, we select equal numbers of correctly and incorrectly classified samples. The remaining samples are used for training. The details of the dataset split are shown in Tab. 1. To address the long-tail problem in the training set, during each training epoch for a mentor, we randomly generate a batch of samples that includes an equal number of correctly and incorrectly classified images by the mentee.

### v2

@@ Abstract
AI models make mistakes when recognizing images—whether in-domain, out-of-domain, or adversarial. Predicting these errors is critical for improving system reliability, reducing costly mistakes, and enabling proactive corrections in real-world applications such as healthcare, finance, and autonomous systems. However, understanding what mistakes AI models make, why they occur, and how to predict them remains an open challenge. Here, we conduct comprehensive empirical evaluations using a “mentor” model—a deep neural network designed to predict another “mentee” model’s errors. Our findings show that the mentor excels at learning from a mentee’s mistakes on adversarial images with small perturbations and generalizes effectively to predict in-domain and out-of-domain errors of the mentee. Additionally, transformer-based mentor models excel at predicting errors across various mentee architectures. Subsequently, we draw insights from these observations and develop an “oracle” mentor model, dubbed SuperMentor, that can outperform baseline mentors in predicting errors across different error types from the ImageNet-1K dataset. Our framework paves the way for future research on anticipating and correcting AI model behaviors, ultimately increasing trust in AI systems. All code, models, and data will be made publicly available.
@@ 1 Introduction
Accurately predicting these errors is critical to enhancing the robustness and reliability of AI, especially in high-stakes real-world applications such as healthcare (Habehh & Gohel, 2021), finance (Mashrur et al., 2020), and autonomous driving (Huang et al., 2022). Proactively identifying potential errors enables more efficient corrections, reducing costly mistakes and safeguarding against catastrophic failures. By predicting when models are likely to err, we can implement strategies that either mitigate or entirely avoid the risks associated with those errors, ultimately leading to more trustworthy AI deployments.
@@
Subsequent research in out-of-domain detection has demonstrated that a model’s accuracy is often correlated with how far the data deviates from in-domain samples (Hendrycks & Dietterich, 2019; Shankar et al., 2021; Li et al., 2017). These methods typically rely on predefined metrics, such as model parameter distances (Yu et al., 2022), model disagreements (Jiang et al., 2021; Madani et al., 2004) and confidence scores (Guillory et al., 2021), which limits their ability to generalize predictions across various data types, including errors arising from in-domain data or adversarial attacks (Szegedy, 2013). In parallel, earlier efforts in trustworthiness prediction mostly depend on shallow neural networks (Corbière et al., 2019; Qiu & Miikkulainen, 2022; Jiang et al., 2018) or carefully-designed loss functions (Luo et al., 2021). However, they do not investigate how different error types influence trustworthiness prediction.
Another line of research improves the robustness of the AI models with adversarial training approaches (Ilyas et al., 2019; Gowal et al., 2020; Balunović & Vechev, 2020); however, these approaches primarily focus on improving the model’s overall performance rather than predicting when errors may occur in the models. Moreover, unlike selective prediction (Geifman & El-Yaniv, 2017), rejection learning (Cortes et al., 2016), and learning to defer (Madras et al., 2018), which jointly train the selection/rejection function with the mentee, our method trains the mentor and mentee independently. This separation is especially beneficial when mentee training is time- and resource-intensive or when its training data is inaccessible.
@@
Different from all these works, we delve into the underlying principles of errors generated by AI models in the task of image classification with another AI model. Specifically, we designate the AI model that predicts errors as the mentor and the AI model being evaluated for performance as the mentee. The mentor strives to predict whether the mentee makes a mistake for any given data. See Fig. 1 for the illustration of the problem setup. Training the mentor on the error patterns made by the mentee can potentially reveal the strengths and weaknesses of the mentee’s learned representations across various visual contexts.
Our main contributions are highlighted below:
@@
1. We conduct an in-depth analysis of how training mentors on each of three distinct error types specified by the mentees—In-Domain (ID) Errors, Out-of-Domain (OOD) Errors, and Adversarial Attack (AA) Errors—affect the performance of error predictions over three increasingly complex image datasets CIFAR-10 (Krizhevsky et al., 2009), CIFAR-100 (Krizhevsky et al., 2009) and ImageNet-1K (Deng et al., 2009). Our results reveal that training mentors with adversarial attack errors from the mentee has the most significant impact on improving the mentor’s error prediction accuracy.
@@
2. We investigate how various mentor model architectures affect error prediction performance. Our experiments demonstrate that transformer-based mentor models outperform other architectures in predicting errors.
@@
3. We explore how varying levels of distortion in OOD and adversarial images affect the accuracy of error predictions. The findings indicate that training mentors on images with small perturbations can improve error prediction accuracy. In addition, we show that a mentor trained to learn error patterns from one mentee can successfully generalize its error predictions to another mentee.
@@ 2 Related work
With the growing deployment of AI models across diverse fields, ensuring their reliability and understanding their limitations has become increasingly crucial. This has led to numerous research in safe AI such as uncertainty estimation (Nado et al., 2021; Lakshminarayanan et al., 2017), anomaly detection (Bogdoll et al., 2022), outlier detection (Boukerche et al., 2020) and out-of-domain detection (Yang et al., 2024). Unlike these areas, which mainly aim to predict whether the input data falls outside the training domain, our focus is on monitoring and predicting errors in AI models by determining whether the model’s output is correct, irrespective of whether the data comes from the training domain. To detect whether the input data is out of scope, the prior approaches mainly rely on softmax outputs (Granese et al., 2021; Hendrycks & Gimpel, 2016; Dang et al., 2024), activations from network layers (Wang et al., 2020; Cheng et al., 2019; Ferreira et al., 2023), shallow neural networks (Corbière et al., 2019; Qiu & Miikkulainen, 2022; Jiang et al., 2018), and carefully-designed loss functions (Luo et al., 2021) in applications such as object detection (Kang et al., 2018) and trajectory prediction (Shao et al., 2023, 2024). However, these methods often rely on manually defined metrics to estimate the likelihood of a mentee making mistakes, or they fail to examine how different error types affect error prediction. In contrast, our strategy employs a separate deep neural network to automatically learn and approximate the mentee’s decision boundaries for specific error types, providing an end-to-end trainable framework for error prediction. Moreover, our research direction differs from selective prediction (Geifman & El-Yaniv, 2017), rejection learning (Cortes et al., 2016), and learning to defer (Madras et al., 2018) by avoiding joint training with the mentee and requiring no knowledge of the mentee’s training data. Besides, although rejecting a mentee’s unreliable predictions and predicting the correctness of a mentee’s outputs are often positively correlated, they remain distinct tasks.
@@
Recent studies (Hendrycks & Dietterich, 2019; Shankar et al., 2021; Li et al., 2017) have shown that a model’s accuracy on a given dataset is often correlated with how far the data deviates from in-domain samples. However, these studies typically rely on pre-defined metrics, such as model parameter distances (Yu et al., 2022), model disagreements (Jiang et al., 2021; Madani et al., 2004), confidence scores (Guillory et al., 2021), domain-invariant representations (Chuang et al., 2020), and domain augmentation (Deng et al., 2021a), limiting their ability to generalize error prediction beyond in-domain data. In contrast, our mentor is capable of predicting both OOD and in-domain errors for a mentee. Additionally, our mentor is an AI model trained end-to-end without relying on manually defined criteria.
@@ 3 Experimental setups
We denote the mentor and mentee networks as and respectively. We also define as the domain-specific set containing all the test images for a mentee, and as their ground-truth object class labels. Therefore, a mentor is expected to make perfect predictions about the correctness of the mentee’s responses (1 for “correct” and 0 for “wrong”) given any image from :
@@ 3.1 Mentors
The first stream generates logits by predicting the probability distribution of a mentee over all the object classes when the mentee classifies the given image. The mentee network is kept fixed while training the mentor. Let us define the mentee’s output logit as . We introduce the distillation loss proposed by (Hinton, 2015): to align with , where represents Kullback-Leibler divergence. is the softmax function. is the temperature, which controls the smoothness of the soft probability distribution.
In the second stream, the mentor is prompted to predict whether the mentee will make a mistake on the given image or not. We denote the predicted binary label as , where 1 indicates that the mentee does not make a mistake and vice versa for 0. This prediction is supervised by , where is the ground truth correctness label of a mentee and represents the mentor’s predicted probability that the mentee’s prediction is correct. The overall loss is .
@@ 3.2 Mentees and their datasets
We employ two architectures as the mentees’ backbones: ResNet50 (He et al., 2016), which is a 2D-Convolutional neural network (2D-CNN), and ViT (Dosovitskiy, 2020), which is a transformer based on self-attention mechanisms.
To train and test our mentees, we include three image datasets of varying sizes and follow their standard data splits: CIFAR-10 (C10, (Krizhevsky et al., 2009)) with 10 object classes, CIFAR-100 with 100 object classes (C100, (Krizhevsky et al., 2009)) and ImageNet-1K with 1000 object classes (IN, (Deng et al., 2009)). The multi-class recognition accuracy on the standard test sets of C10, C100 and IN datasets are 96.98%, 84.54%, 76.13% for the ResNet50 mentee and 97.45%, 86.51%, 81.07% for the ViT mentee respectively. The parameters of the mentees are frozen in all the experiments conducted on mentors.
@@ 3.3 Datasets for training and testing mentors
The mentor’s objective is to predict whether the mentee will misclassify a given image, regardless of its source. The mentor is trained on correctly and wrongly classified images by a mentee. Next, we introduce how these images are curated and collected. A mentee may encounter various types of errors when dealing with real-world data. To explore which error types most effectively reveal the mentee’s learning patterns, we categorize errors into three types: (1) errors from in-domain images, (2) errors from out-of-domain images, and (3) errors from adversarial images generated by adversarial attacks. Next, we introduce these three error types in detail.
In-Domain (ID) Errors occur on data that come from the same domain as the mentee’s training dataset. Specifically, errors on images from the standard validation set of ImageNet-1K or the test sets of CIFAR-10 and CIFAR-100 are considered ID errors. Along with the correctly classified images from these standard test sets, we create three datasets for a mentor: IN-ID, C10-ID, and C100-ID, following the naming convention of [Dataset]-[Error Type].
Out-of-domain (OOD) Errors refer to errors that arise when the mentee encounters data outside the training domain. To obtain OOD samples of a dataset, we adopt four types of image corruptions from (Hendrycks & Dietterich, 2019): speckle noise (SpN) (noise category), Gaussian blur (GaB) (blur category), spatter (Spat) (weather category), and saturate (Sat) (digital category). The noise levels can vary and we select level 1 for image corruptions as specified in (Hendrycks & Dietterich, 2019) by default. As noise levels increase, distortions on OOD images become more pronounced, causing the mentee to make more errors.
@@
Note that adversarial attacks are not always successful, and mentees can still correctly classify some adversarial images. We collect both the correctly and incorrectly classified adversarial images by a mentee based on C10 images, curating four datasets for the mentor: C10-AA-PGD, C10-AA-CW, C10-AA-Jitter, C10-AA-PIFGSM. Similarly, we can also curate four datasets each for a mentor based on C100 and IN images of a mentee.
@@
Training and test splits. The number of correctly and incorrectly classified images by a mentee can vary significantly based on its classification performance. To create balanced training and test sets for the mentor and avoid the long-tailed distribution problem, we select an equal number of correctly and incorrectly classified samples for each training batch and every test set. See Appendix, Sec. A for more details.
3.4 Baselines and evaluation metric
Baselines. We include seven baseline methods for error prediction of a mentee. 1) Self Error Rate (SER) predicts the correctness of the mentee’s outputs by randomly assigning “correct” or “wrong” based on the mentee’s in-domain accuracy; 2) Maximum Class Probability (MCP) (Hendrycks & Gimpel, 2016) classify the mentee’s predictions as correct if its MCP exceeds a predefined threshold. 3) Class Probability Entropy (CPE) evaluates a mentee’s prediction as correct if its entropy of the probability distribution over all classes is below a predefined threshold. 4) Distance To Centroid (DTC) considers a prediction correct if its feature embedding is within a predefined distance from the class centroid. 5) ConfidNet (Corbière et al., 2019) predicts failure by learning True Class Probability (TCP) on the training samples. 6) TrustScore (Jiang et al., 2018) employs Trust Score to quantify the agreement between the classifier and a modified nearest-neighbor classifier. 7) Steep Slope Loss (SSL) (Luo et al., 2021) develops a neural network for predicting trustworthiness by utilizing a steep slope loss function. See Appendix, Sec. B for more details.
Evaluation metric. To assess the performance of mentors, we report their error prediction accuracy on the test set corresponding to each specified error source. For instance, a mentor trained on the C10-ID training set is evaluated on the C10-OOD-SpN test set. The error prediction accuracy is calculated by averaging the mentor’s accuracies on the samples that the mentee correctly classified and those that the mentee incorrectly classified. However, since a mentee can make mistakes across various real-world scenarios, a mentor must accurately predict errors across all error types. Therefore, we compute the average accuracy, named as Accuracy (%) of a mentor across all test sets, including one ID error, four OOD errors, and four AA errors. See Appendix, Sec. C for an example calculation of the average accuracy of a mentor.
4 Results
A mentee’s mistakes can reveal their learning biases, behaviors, or traits. Here, we investigate which types of errors offer the most insight into understanding a mentee’s decision boundaries during image recognition tasks. We train mentors with identical architectures on datasets containing specific error types made by the mentee across C10 (Fig. 3(a)), C100 (Fig. 3(b)), and IN (Fig. 3(c)). For instance, if a mentor trained on C10-OOD achieves higher accuracy in error prediction compared to one trained on C10-IN, this suggests that in-domain errors provide less diagnostic information about the mentee’s decision-making process than out-of-domain errors. Both mentors and mentees may have the same or different backbones, such as ResNet50 (R) or ViT (V).
As shown in Fig. 3, over C10, C100, and IN images, the high accuracy for mentors trained on adversarial attack (AA) errors indicates that these errors offer deeper insights into the mentee’s decision process compared to out-of-domain (OOD) and in-domain (ID) errors. In some cases, mentors trained on OOD errors slightly outperformed those trained on ID errors, though both were still inferior to those trained on AA errors. Interestingly, our experiments reveal that training mentors exclusively on AA errors performs marginally worse than joint training on ID, OOD, and AA errors (see Appendix, Sec. D). This result underscores a significant overlap of mentee error patterns among ID, OOD, and AA data.
Loss landscape analysis. A loss landscape of a mentee reflects how a mentee’s loss function behaves across different parameter configurations. Mentors’ performance offers insights into the structure of a mentee’s loss landscape. Consistent with (Ilyas et al., 2019), the high accuracy of mentors trained on AA errors suggests that adversarial images lie closer to the mentee’s decision boundary, enabling more accurate prediction of the mentee’s mistakes and a deeper understanding of the loss landscape. Similarly, OOD data aids mentors in learning boundaries by shifting ID samples closer to the boundary. However, it does not explore the boundary as thoroughly as adversarial images. ID data, with fewer samples near the boundary, provides more limited exploration compared to adversarial images.
4.2 Mentor architectures matter in error predictions
Loss landscape analysis. The performance difference between mentors’ architectures is due to ViT’s superior ability to identify features from error patterns. Its self-attention mechanism captures complex relationships among data samples, providing a deeper understanding of the mentee’s loss landscape, particularly in modelling irregular, rugged landscapes with sharp peaks and valleys.
| ID | SpN | GaB | Spat | Sat | PGD | CW | Jitter | PIFGSM | Average | |
| SER | 50.2 | 49.8 | 49.9 | 50.1 | 49.9 | 50.1 | 49.9 | 49.9 | 49.9 | 50.0 |
| MCP () | 77.3 | 76.8 | 77.4 | 78.3 | 76.6 | 71.8 | 72.9 | 71.7 | 69.4 | 74.7 |
| CPE () | 78.0 | 75.2 | 75.9 | 78.0 | 76.2 | 72.2 | 72.8 | 72.7 | 69.3 | 74.5 |
| DTC () | 52.4 | 51.2 | 52.9 | 52.7 | 52.5 | 51.4 | 51.6 | 51.7 | 51.6 | 52.0 |
| ConfidNet (Corbière et al., 2019) | 67.3 | 65.9 | 68.5 | 70.8 | 67.9 | 68.8 | 69.0 | 70.1 | 67.1 | 68.4 |
| TrustScore (Jiang et al., 2018) | 70.0 | 67.5 | 71.1 | 73.4 | 71.3 | 73.2 | 72.5 | 73.9 | 72.7 | 71.7 |
| SSL (Luo et al., 2021) | 66.3 | 66.7 | 69.7 | 71.9 | 70.0 | 69.1 | 70.1 | 70.2 | 66.7 | 69.0 |
| SuperMentor (ours) | 78.9 | 73.6 | 74.6 | 78.3 | 73.6 | 83.0 | 78.4 | 79.9 | 72.3 | 77.0 |
Although adversarial images have been demonstrated to aid in error prediction (Sec. 4.1), it remains unclear whether adversarial images with varying degrees of image distortion exhibit the same effect. A straightforward method to regulate the level of image distortion caused by adversarial attacks is to set the perturbation bound . We employ four corruption levels by setting , , , and . We use the adversarial attack PIFGSM as an example since the error patterns from PIFGSM are most effective for the mentor’s prediction (see Appendix, Fig. S3- S5). As shown in Fig. 5, the mentor’s accuracy significantly decreases as the distortion level increases. In particular, for the C10-AA-PIFGSM, the accuracy at level 1 is , which is notably higher than at level 4. Our findings suggest that adversarial attacks employing smaller perturbations yield more benefits for mentor error prediction. This phenomenon can be attributed to the fact that adversarial images with minimal perturbations maintain closer proximity to the decision boundary of a mentee.
4.4 Mentors generalize across mentees
In Sec. 4.1, mentors have demonstrated their ability to learn the error patterns of mentees. This observation raises an important question: can the error patterns learned from one mentee (mentee A) be generalized to another mentee (mentee B) when the two mentees employ different model architectures? To explore this, we evaluate all 324 mentors, whose performances are depicted in Appendix, Tab. S6- S8, on the alternate mentee. Specifically, mentors trained on the errors of the ResNet50 mentee are tested on the predictions of the ViT mentee, and vice versa. The outcomes of these evaluations are illustrated in Fig. 5. Surprisingly, most points lie near the dashed diagonal line, implying that the mentors’ performance does not significantly deteriorate when evaluated on the predictions of different mentee architectures. This finding indicates that ResNet50 and ViT mentees tend to produce similar error patterns when trained on the same dataset.
4.5 Our SuperMentor outperforms all baselines
By drawing insights from observations in the subsections above, we propose an “oracle” mentor model, dubbed SuperMentor. We introduce the technical novelties of our SuperMentor below. First, as demonstrated in Sec. 4.1 and Sec. 4.3, mentors trained on adversarial images with small perturbations of a mentee outperform those trained on OOD and ID images; thus, our SuperMentor adopts the training data from the PIFGSM error source of mentees with . Second, since ViT has been proven to be a more effective architecture for mentors than ResNet50 (Sec. 4.2), SuperMentor adopts ViT as the backbone architecture.
We demonstrate the effectiveness of SuperMentor by comparing it with baselines introduced in Sec. 3.4. The results are shown in Tab. 1. It can be seen that our SuperMentor consistently achieves higher average accuracies over all error types compared to all baselines. Notably, in the AA scenarios, SuperMentor has significantly higher accuracy than the baselines. For example, SuperMentor achieves an error prediction accuracy of 83.0% on IN-AA-PGD of the ResNet50 mentee, whereas the best baseline only reaches 73.3%.
Baseline methods like SER, MCP, CPE, and DTC rely on fixed thresholds or predetermined values of manually defined criteria such as confidence or entropy, making them less adaptable to OOD and AA scenarios. In contrast, SuperMentor leverages deep neural networks to capture the complexity of error prediction. Additionally, methods like ConfidNet, TrustScore, and SSL are trained on ID error types. As shown in Tab. 1, these baselines are inferior to SuperMentor, which is trained exclusively on AA data. This suggests that the source of error data used to train the mentor plays a more critical role than the choice of loss functions or the sophisticated training strategies in these baselines.
We conduct extra network analysis of our SuperMentor. Specifically, we include an ablation study on the distillation loss (Appendix, Sec. F). The removal of leads to a significant drop in Accuracy, highlighting the importance of . Moreover, we provide the visualization of SuperMentor’s embeddings based on the correctness of a mentee’s prediction across three error types (Appendix, Sec. G). Two distinct clusters are observed between samples wrongly or correctly classified by the mentee. This suggests that our SuperMentor is capable of accurately predicting the errors of a mentee regardless of the error sources.
5 Discussion and Conclusion
In our work, we tackle the challenge of predicting errors of AI models through extensive empirical evaluations using an end-to-end trainable “mentor” model. This mentor is designed to assess the correctness of a mentee’s predictions across three distinct error types: in-domain errors, out-of-domain errors, and adversarial attack errors. Our results show that the mentor excels at learning from a mentee’s errors on adversarial images with minimal perturbations and, surprisingly, generalizes well to both in-domain and out-of-domain predictions of the same mentee. Additionally, we highlight the effectiveness of transformer-based mentor architectures compared to 2D-CNN-based ones, demonstrating their superior generalization capabilities across mentees with diverse backbones. Lastly, we introduce the SuperMentor, which outperforms all existing mentor baselines.
Despite the promising results, our framework may raise concerns regarding the vulnerabilities of AI mentors. Indeed, vulnerability is a well-known challenge for all Deep Neural Network (DNN)-based systems, yet this has not hindered their transformative applications in the real world. For example, DNNs are used to detect diabetic retinopathy (Jumper et al., 2021; Alyoubi et al., 2020), outperforming traditional diagnostic methods. Similarly, our work uses one AI to diagnose the other AI, which is a valid and impactful use case. To demonstrate its real-world utility, we extended our experiments to a medical image classification task for colorectal cancer diagnosis (Appendix, Sec. H). Our mentor achieves high error prediction accuracy compared to all competitive baselines, showcasing its practical uses in high-stake applications. More importantly, beyond the OOD domains evaluated in Sec. 4.5, results in Appendix, Sec. I show that our SuperMentor remains robust to mentee error patterns across an additional set of five unseen OOD domains. Moreover, the vulnerabilities of the mentor AI do not overlap with those of the mentees since mentors operate independently of the mentee’s architecture or training details, reducing the risk of shared weaknesses. To validate this point, we use EigenCAM to visualize the behavior of both our ViT SuperMentor and ViT mentee on example images (Appendix, Fig. S2). From Fig. S2 (a), the mentor does not merely replicate the mentee’s learning patterns, as their activation maps for identical input images can differ significantly. Interestingly, our AI mentor can even serve as a valuable diagnostic tool for the AI mentees. In Appendix, Fig. S2 (b), the mentor highlights the key feature pattern diagnostic for object recognition that the mentee tends to overlook. This suggests that the mentor can identify vulnerabilities in DNN-based models.
Our work paves the way for several promising research directions in safe and trustworthy AI. First, while our current research focuses on image classification, there is potential to extend this approach to other vision and language tasks, such as object detection and machine translation. Second, future research could explore mutual learning between mentors and mentees, where mentors not only learn from the mentee’s error patterns but also provide valuable feedback to help refine the mentee. Third, we can establish more rigorous evaluation criteria for mentors, broadening their predictive capabilities. For example, beyond predicting whether a mentee is likely to make errors, mentors could also forecast the specific types of errors a mentee may encounter. Fourth, drawing parallels with AI mentors for AI mentees, we can explore the possibility of using AI mentors to investigate recognition errors in humans and primates. Such studies could provide insights into error pattern alignment between biological and artificial intelligent systems.
Impact Statement
As AI integrates into our daily lives and critical applications like healthcare, finance, and autonomous driving, ensuring its safety and reliability is imperative. Our research highlights the error patterns and characteristics of AI models for object recognition, paving the way for building trustworthy AI. By enabling one AI model to anticipate and correct another’s behavior, our framework helps ensure AI systems operate in a more predictable and reliable manner. Overall, our work lays the foundation for developing systems capable of anticipating the errors of others, offering practical value in high-stakes real-world applications.
Appendix A Dataset split for each error source used in mentor training
As mentioned in Sec. 3.3, for any given dataset of a mentor, let and represent the sets of correctly and incorrectly classified images by a mentee. The sizes of and can vary significantly, depending on the mentee’s classification performance. A mentee with high recognition accuracy will have more correct classifications (big ) and fewer incorrect ones (small ). To create a balanced test set for a mentor, we select equal numbers of correctly and incorrectly classified samples. The remaining samples are used for training. The details of the dataset split are shown in Appendix, Tab. S1. To address the long-tail problem in the training set, during each training epoch for a mentor, we randomly generate a batch of samples that includes an equal number of correctly and incorrectly classified images by the mentee.